OpenStudy (anonymous):

1 over 3 (2x − 8) = 4

OpenStudy (karatechopper):

well..start off by writing it numerically!

OpenStudy (anonymous):

I don't know how to do that..

OpenStudy (anonymous):

2/3x-8/3=4 2/3x=20/3 x=10